please help me find the slope of the ff:
a. (9,8) (3,0)
b. (-4,-4) (4,-9)
thank you.

Answer by Mike(39)

The formula to find the slope which you should use for these types of questions is:
(y2-y1)/(x2-x1)
where y2 is the second y coordinate and y1 is the first y coordinate (the same rule applies for x2 and x1).
Thus,
a) (0-8)/(3-9) = 8/6 = 4/3
b) (-9+4)/(4+4) = -5/8
Hope that helps.
Regards,
-Mike